# FIHA: Autonomous Hallucination Evaluation in Vision-Language Models with Davidson Scene Graphs

### v1

@@ Abstract
The rapid development of Large Vision-Language Models (LVLMs) often comes with widespread hallucination issues, making cost-effective and comprehensive assessments increasingly vital. Current approaches mainly rely on costly annotations and are not comprehensive – in terms of evaluating all aspects such as relations, attributes, and dependencies between aspects. Therefore, we introduce the FIHA (autonomous Fine-graIned Hallucination evAluation evaluation in LVLMs), which could access hallucination LVLMs in the LLM-free and annotation-free way and model the dependency between different types of hallucinations. FIHA can generate Q&A pairs on any image dataset at minimal cost, enabling hallucination assessment from both image and caption. Based on this approach, we introduce a benchmark called FIHA-v1, which consists of diverse questions on various images from MSCOCO and Foggy. Furthermore, we use the Davidson Scene Graph (DSG) to organize the structure among Q&A pairs, in which we can increase the reliability of the evaluation. We evaluate representative models using FIHA-v1, highlighting their limitations and challenges. We released our code and data [^1] https://anonymous.4open.science/r/FIHA-45BB.
@@ 1 Introduction
Evaluation Methods Discriminative Hallucination Task Type Use DSG LLM Free Annotation Free
@@
POPE [26] $\times$ $\times$ $\times$ $\times$
NOPE [31] $\times$ $\times$ $\times$ $\times$ $\times$
CIEM [15] $\times$ $\times$ $\times$ $\times$
Bingo [8] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
AMBER [37] $\times$ $\times$
HallusionBench [28] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
MHaluBench [4] $\times$ $\times$ $\times$ $\times$ $\times$
Hal-Eavl [17] $\times$ $\times$
@@
Large Vision-Language Models (LVLMs) such as MiniGPT-4  [45] and LLaVA  [30], which extend Large Language Models (LLMs) by incorporating visual encoders, such as CLIP [34], have shown prominent capabilities in visual understanding and generation [42]. However, LVLMs suffer from the issue of hallucination, which can lead to misinterpretation or erroneous assertions of the visual inputs, thus hindering the performance of models in multi-modal tasks [16, 21]. Specifically, the models may describe objects that do not exist in the image or incorrect object attributes and relations between objects. Generating such unreliable content will greatly reduce the model’s credibility. Therefore, it is crucial to establish a benchmark for evaluating the hallucination level of LVLMs.
Previous studies [26, 38, 37], as shown in Table 1, primarily employ a Question Generation (QG) module to create a set of validation questions and expected answers (i.e. Q&A pairs) for hallucination evaluation. These generated questions are then used to evaluate hallucinations in LVLMs. Despite the compelling success of the existing work, they still face two main challenges: (1) The existing work overlooks the dependency between different kinds of questions. For example, if the answer to “Is there a bike?” is no, dependent questions like “Is the bike yellow?” should be skipped. (2) Additionally, most prior work heavily relies on human annotations [37] or LLMs [24] to generate Q&A pairs used in hallucination evaluation, which can be costly or labor-intensive.
To mitigate these limitations, we propose Fine-grained Hallucination Evaluation (FIHA), an automatic evaluation framework for assessing fine-grained and diverse hallucinations in large-scale vision-language models. The framework accepts either images or captions as input and generates Q&A pairs by extracting objects, attributes, and entity relations. It then formulates diverse questions (e.g., “what”, “who”, “which”, etc.) that allow for free-form responses. By integrating BLIP-2 [25] for caption generation, Fast R-CNN [11] for feature extraction, and a question-generation template, our pipeline enables fully automatic Q&A generation without relying on LLMs or manual annotations.
To organize all the Q&A pairs into a tree-like structure, we introduce the Davidson Scene Graph (DSG) [5]. The DSG ensures that responses to leaf node questions depend on the correctness of the root node answers. This structure mitigates situations where a large language model (LLM) may provide correct answers to fine-grained questions while hallucinating basic attributes. By enforcing this dependency, we aim to reduce errors and improve the reliability of the model’s responses. In addition, our Q&A pairs cover various types of questions, including negative, narrative, and interrogative questions. This type of tree structure allows for a progressive deepening of questions, enabling a comprehensive evaluation of the model’s understanding of the image.
@@ 2 Methodology
Figure 1: Overview of FIHA framework. FIHA extracts entities, attributes, and relations from images and captions respectively, and generates comprehensive and diverse questions to thoroughly detect model hallucinations. In the Figure, we can see that no large language model (LLM) [1] or additional manual annotations are used.
In this section, we introduce the overall pipeline of FIHA as illustrated in Figure 1. In summary, our pipeline offers two approaches for Q&A generation. The first is based on images: given an image $I$, we extract the necessary entities, including features such as objects, object attributes, and relations. Using a rule-based method, we then generate the Q&A pairs. The second approach is based on captions. If an image does not already have a caption, we can use BLIP-2 [25] to generate captions. Alternatively, if the dataset includes original captions, we can use them directly as input, pass them through the feature extraction process, and generate the corresponding Q&A pairs.
@@ 2.1.1 Information Acquisition from the Caption
Caption Generation. Image captions can depict an image in detail, demonstrating fine-grained visual information, such as objects, attributes and relations. Inspired by the findings of previous research [26], which indicate that smaller vision-language models tend to produce more concise responses with fewer hallucinations compared to mainstream LVLMs. As such, we select BLIP-2 to generate a caption for the image. This step allows us to generate highly credible captions based on the image.
Fine-grained Information Extraction. In this process, we take either the generated captions (if the ground-truth caption is not available) or the ground-truth captions, depending on the user’s needs, as input and extract information such as object existence, object attributes, and relations from the captions. For extracting objects and attributes in the caption, we use SpaCy’s [14] part-of-speech tagging feature to identify objects and their corresponding attributes, such as numerals, adjectives, and verbs. As a result, we obtain all the ground truth objects and their attributes as: $G^{C}_{O,A}=\{o_{1}:A_{1},o_{2}:A_{2},\dots,o_{n}:A_{n}\}$, where $n$ is the number of objects. $o_{i}$ is the $i$-th object and $A_{i}$ is all attributes for the $i$-th object. Relations from the captions are extracted using the Stanford CoreNLP library [^2] https://stanfordnlp.github.io/CoreNLP/., which provides a powerful suite of NLP tools for performing various linguistic analyses on text, making it an ideal choice for relation extraction. From this process, we obtain all the relations: $G^{C}_{R}=\{R_{1}(o_{R_{1}}^{1},o_{R_{1}}^{2}),R_{2}(o_{R_{2}}^{1},o_{R_{2}}^{2}),\dots,R_{m}((o_{R_{m}}^{1},o_{R_{m}}^{2}))\}$, where $m$ is the number of relations. $R_{i}$ is the $i$-th relation for the objects $o_{R_{i}}^{2}$ and $o_{R_{i}}^{2}$.
@@ 2.1.2 Information Acquisition from the Image
As the caption may lose some information in the image, our second approach to extract information is directly from images, which provides richer and more detailed information than captions alone. For object and attribute extraction, we use Fast R-CNN [11], a well-established and widely used object detection method based on Region-based Convolutional Networks. Fast R-CNN has been a pioneering approach in the field of object detection due to its ability to quickly identify objects within an image while simultaneously predicting their attributes. This method allows us to retrieve the ground truth objects along with their corresponding attributes such as color, size, and shape, forming a set of objects and attributes: $G^{I}_{O,A}=\{o_{1}:A_{1},o_{2}:A_{2},\dots,o_{n}:A_{n}\}$, where $n$ represents the number of objects detected. In addition to identifying objects and their attributes, it is crucial to understand how these objects interact or relate within the scene. For this purpose, we employ RelTR [6], a cutting-edge method designed to generate sparse scene graphs by decoding visual appearances and learning both subject and object queries from the image data. RelTR enables us to extract meaningful relationships between the detected objects, such as spatial relations (e.g., one object being behind or near another) and actions (e.g., wearing or holding), resulting in a set of ground truth relations: $G^{I}_{R}=\{R_{1}(o_{R_{1}}^{1},o_{R_{1}}^{2}),R_{2}(o_{R_{2}}^{1},o_{R_{2}}^{2}),\dots,R_{m}((o_{R_{m}}^{1},o_{R_{m}}^{2}))\}$, where $m$ denotes the number of relationships extracted from the image. By combining both object detection and relational extraction, this approach provides a comprehensive understanding of the visual content, which is essential for generating accurate and meaningful Q&A pairs.
@@ 2.2 Question Answer Pair Generation
We use two kinds of questions for hallucination evaluation. The first type is Yes-No questions, which check object existence, such as “Is there any {objk}?” and object relations like “Is there a {obj1} near the {objk}?”. These questions help determine whether specific objects and their relationships are present within an image. Additionally, Yes-No questions assess attributes by asking about features like color, size, or location.
The second type is Wh-Questions, which add diversity to the evaluation by incorporating interrogative words such as “what”, “who”, “which”, “where”, and “how many”. These questions elicit more detailed, free-form responses, typically no longer than three words. For example, “What color is the {objk}?” or “Which object is near the {objk}?” help assess the finer details about objects and their relationships. Unlike traditional hallucination evaluations that primarily rely on Yes-No questions, our approach includes both types to provide a more comprehensive assessment.
Additionally, we introduce Negative Questions for both Yes-No questions and Wh-Questions. These questions are generated by replacing the objects, their attributes, and relations in the original Q&A pairs with non-existent ones. Specifically, we randomly select several pieces of information from a complete list of objects, attributes (such as color, size, etc.), and relations to replace the real information. Such questions are answered with negative pronouns such as “none”, “nobody”, or “nowhere”.
@@ 2.3 Davidson Scene Graph
To model the dependency between objects, attributions, and relations accurately and improve the reliability of hallucination evaluation, we introduce the Davidsonian Scene Graph (DSG) [5] mechanism. The DSG can be understood as a post-processing step for the Q&A pairs. After obtaining all the Q&A pairs, we organize them into multiple tree-like structures, where each Q&A pair serves as a node. According to the structure of the tree-like structures, each node is either a root node or a leaf node. Specifically, the entire process is divided into three steps. In step 1, we set the question about the existence of a certain object as the root node. In step 2, we set all questions related to the object of the root node, such as those about its attributes and relations, as corresponding leaf nodes. Finally, in step 3, determine whether the root node question is answered correctly; if not, there is no need to judge the questions at the leaf nodes, and we directly determine that all questions on the tree are answered incorrectly. For instance, after step 1 and step 2, we obtain a list of questions such as $L^{Q}=\{Q_{1}:Independent,Q_{2}:Depends\ on\ Q_{1}\}$. Before determining if the answer to $Q_{2}$ is correct, we first assess $Q_{1}$, which concerns the accuracy related to the root node. If the question about the existence of an object, which is at the root node, is answered incorrectly, we consider that all other related questions must be hallucinatory.
@@ 3 Experiments
Table 2: Data statistics of FIHA-v1. #Num denotes the number of the images.
Source #Num From Image From Caption
MSCOCO 500 25,699 13,007
Foggy 150 7,232 2,801
Visual Genome 50 1,566 476
Table 3: Evaluation results of LVLMs on questions generated from images and captions using FIHA. The upper part is from the MSCOCO dataset and the bottom part is from the Foggy dataset.
@@
mPLUG-Owl 42.1 70.2 61.4 43.7 31.4 61.6 55.5 31.2
MiniGPT-4 23.5 27.5 22.2 22.1 15.9 25.7 28.8 14.2
MultiModal-GPT 59.1 46.4 47.1 46.6 23.8 39.6 45.7 22.1
LLaVA-1.5-7B 77.8 77.0 65.9 67.7 50.7 64.9 67.5 50.5
LLaVA-1.5-13B 78.9 80.9 66.4 68.3 47.6 64.2 65.5 48.5
InstructBLIP 84.7 83.3 78.6 80.4 65.7 69.5 77.4 64.2
GPT-4V 87.2 81.4 86.3 85.5 70.3 71.5 75.8 69.3
@@
mPLUG-Owl 64.8 60.2 51.1 42.7 29.5 58.9 51.6 25.6
MiniGPT-4 30.1 30.2 27.6 28.1 23.4 34.4 37.8 23.0
MultiModal-GPT 50.2 48.7 46.1 45.8 28.1 43.9 47.9 25.4
LLaVA-1.5-7B 67.7 68.4 56.2 52.9 29.1 50.0 49.2 25.8
LLaVA-1.5-13B 68.1 71.5 56.1 52.3 28.9 49.2 49.8 25.5
InstructBLIP 70.9 75.6 60.2 58.8 32.8 58.3 53.2 30.5
GPT-4V 76.3 70.1 64.6 66.0 33.7 53.3 51.7 32.1
@@
mPLUG-Owl 41.8 68.9 60.9 43.3 44.6 71.7 51.8 33.8
MiniGPT-4 22.9 26.8 22.0 21.8 15.9 25.7 28.8 14.2
MultiModal-GPT 58.8 46.1 46.9 46.3 65.3 65.2 62.2 61.7
LLaVA-1.5-7B 77.7 77.2 61.1 67.9 56.4 73.8 62.0 52.6
LLaVA-1.5-13B 79.0 81.2 66.7 68.6 74.3 79.6 68.8 69.2
InstructBLIP 84.5 83.7 79.0 80.7 67.7 78.4 71.9 66.7
GPT-4V 87.0 81.2 86.0 85.3 84.2 78.9 84.1 82.2

### v2

# FIHA: Automated Fine-grained Hallucinations Evaluations in Large Vision Language Models with Davidson Scene Graphs

@@ Abstract
The rapid development of Large Vision-Language Models (LVLMs) often comes with widespread hallucination issues, making cost-effective and comprehensive assessments increasingly vital. Current approaches mainly rely on costly annotations and are not comprehensive – in terms of evaluating all aspects, such as relations, attributes, and dependencies between aspects. Therefore, we introduce the FIHA (autonomous Fine-graIned Hallucination evAluation in LVLMs), which could access LVLMs hallucination in an LLM-free and annotation-free way and model the dependency between different types of hallucinations. FIHA can generate Q&A pairs on any image dataset at minimal cost, enabling hallucination assessment from both image and caption. Based on this approach, we introduce a benchmark called FIHA-v1, which consists of diverse questions on various images from three datasets. Furthermore, we use the Davidson Scene Graph (DSG) to organize the structure among Q&A pairs, in which we can increase the reliability of the evaluation. We evaluate representative models using FIHA-v1, highlighting their limitations and challenges. We released our code and data at https://github.com/confidentzzzs/FIHA.
@@ 1 Introduction
Evaluation Methods Discriminative Hallucination Task Type Dependency LLM Free Annotation Free
@@
POPE Li et al. (2023d) $\times$ $\times$ $\times$ $\times$
NOPE Lovenia et al. (2023) $\times$ $\times$ $\times$ $\times$ $\times$
CIEM Hu et al. (2023) $\times$ $\times$ $\times$ $\times$
Bingo Cui et al. (2023a) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
AMBER Wang et al. (2023a) $\times$ $\times$
HallusionBench Liu et al. (2023a) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
MHaluBench Chen et al. (2024) $\times$ $\times$ $\times$ $\times$ $\times$
Hal-Eavl Jiang et al. (2024a) $\times$ $\times$
@@
Large Vision-Language Models (LVLMs) such as MiniGPT-4  Zhu et al. (2023) and LLaVA  Liu et al. (2023c), which extend Large Language Models (LLMs) by incorporating visual encoders, such as CLIP Radford et al. (2021), have shown prominent capabilities in visual understanding and generation Zhang et al. (2024). However, LVLMs suffer from the issue of hallucination, which can lead to misinterpretation or erroneous assertions of the visual inputs, thus hindering the performance of models in multi-modal tasks Huang et al. (2023); Jing et al. (2024); Zhang et al. (2025). Specifically, the models may describe objects that do not exist in the image or incorrect object attributes and relations between objects. Generating such unreliable content will greatly reduce the model’s credibility. Therefore, it is crucial to establish a benchmark for evaluating the hallucination level of LVLMs.
Previous studies Li et al. (2023d); Wang et al. (2023b, a), as shown in Table 1, primarily employ a Question Generation (QG) module to create a set of validation questions and expected answers (i.e., Q&A pairs) for hallucination evaluation. These generated questions are then used to evaluate hallucinations in LVLMs. Despite the compelling success of the existing work, they still face two main challenges: (1) The existing work overlooks the dependency between different kinds of questions. For example, if the answer to “Is there a bike?” is no, dependent questions like “Is the bike yellow?” should be skipped, detailed explanations can be found in the Appendix C. (2) Additionally, most prior work heavily relies on human annotations Wang et al. (2023a) or LLMs Li et al. (2023c) to generate Q&A pairs used in hallucination evaluation, which can be costly or labor-intensive.
To mitigate these limitations, we propose Fine-grained Hallucination Evaluation (FIHA), an automatic evaluation framework for assessing fine-grained and diverse hallucinations in large-scale vision-language models. The framework accepts either images or captions as input and generates Q&A pairs by extracting objects, attributes, and entity relations. It then formulates diverse questions (e.g., “what”, “who”, “which”, etc.) that allow for free-form responses. By integrating BLIP-2 Li et al. (2023b) for caption generation, Fast R-CNN Girshick (2015) for feature extraction, and a question-generation template, our pipeline enables fully automatic Q&A generation without relying on LLMs or manual annotations.
We organize Q&A pairs using the Davidson Scene Graph (DSG) Cho et al. (2023). The DSG ensures leaf node responses depend on root node answers, reducing errors and improving reliability. Our Q&A pairs include negative, narrative, and interrogative questions, allowing a progressive, comprehensive evaluation of image understanding.
@@ 2 Methodology
Figure 1: Overview of FIHA framework. FIHA extracts entities, attributes, and relations from images and captions respectively, and generates comprehensive and diverse questions to thoroughly detect hallucinations. In the Figure, we can see that no large language model (LLM) Achiam et al. (2023) or additional manual annotations are used.
In this section, we introduce the overall pipeline of FIHA as illustrated in Figure 1. In summary, our pipeline offers two approaches for Q&A generation. The first is based on images: given an image $I$, we extract the necessary entities, including features such as objects, object attributes, and relations. Using a rule-based method, we then generate the Q&A pairs. The second approach is based on captions. If an image does not already have a caption, we can use BLIP-2 Li et al. (2023b) to generate captions. Alternatively, if the dataset includes original captions, we can use them directly as input, pass them through the feature extraction process, and generate the corresponding Q&A pairs.
@@ 2.1.1 Information Acquisition from the Caption
Caption Generation. Image captions can depict an image in detail, demonstrating fine-grained visual information, such as objects, attributes and relations. Inspired by the findings of previous research Li et al. (2023d), which indicate that smaller vision-language models tend to produce more concise responses with fewer hallucinations compared to mainstream LVLMs. As such, we select BLIP-2 to generate a caption for the image. This step allows us to generate highly credible captions based on the image.
Fine-grained Information Extraction. In this process, we take either the generated captions (if the ground-truth caption is not available) or the ground-truth captions, depending on the user’s needs, as input and extract information such as object existence, object attributes, and relations from the captions. For extracting objects and attributes in the caption, we use SpaCy’s Honnibal and Montani (2017) part-of-speech tagging feature to identify objects and their corresponding attributes, such as numerals, adjectives, and verbs. As a result, we obtain all the ground truth objects and their attributes as: $G^{C}_{O,A}=\{o_{1}:A_{1},o_{2}:A_{2},\dots,o_{n}:A_{n}\}$, where $n$ is the number of objects. $o_{i}$ is the $i$-th object and $A_{i}$ is all attributes for the $i$-th object. Relations from the captions are extracted using the Stanford CoreNLP library [^1] https://stanfordnlp.github.io/CoreNLP/., which provides a powerful suite of NLP tools for performing various linguistic analyses on text, making it an ideal choice for relation extraction. From this process, we obtain all the relations: $G^{C}_{R}=\{R_{1}(o_{R_{1}}^{1},o_{R_{1}}^{2}),\dots,R_{m}((o_{R_{m}}^{1},o_{R_{m}}^{2}))\}$, where $m$ is the number of relations. $R_{i}$ is the $i$-th relation for the objects $o_{R_{i}}^{2}$ and $o_{R_{i}}^{2}$.
@@ 2.1.2 Information Acquisition from the Image
As the caption may lose some information in the image, our second approach to extract information is directly from images, which provides richer and more detailed information than captions alone. For object and attribute extraction, we use Grounding DINO Liu et al. (2023d), a well-established and widely used object detection method based on Transformer architecture. Grounding DINO has been a pioneering approach in the field of object detection due to its ability to quickly identify objects within an image while simultaneously predicting their attributes. This method allows us to retrieve the ground truth objects along with their corresponding attributes such as color, size, and shape, forming a set of objects and attributes: $G^{I}_{O,A}=\{o_{1}:A_{1},o_{2}:A_{2},\dots,o_{n}:A_{n}\}$, where $n$ represents the number of objects detected. In addition to identifying objects and their attributes, it is crucial to understand how these objects interact or relate within the scene. For this purpose, we employ RelTR Cong et al. (2022), a cutting-edge method designed to generate sparse scene graphs by decoding visual appearances and learning both subject and object queries from the image data. RelTR enables us to extract meaningful relationships between the detected objects, such as spatial relations (e.g., one object being behind or near another) and actions (e.g., wearing or holding), resulting in a set of ground truth relations: $G^{I}_{R}=\{R_{1}(o_{R_{1}}^{1},o_{R_{1}}^{2}),\dots,R_{m}((o_{R_{m}}^{1},o_{R_{m}}^{2}))\}$, where $m$ denotes the number of relationships extracted from the image. By combining both object detection and relational extraction, this approach provides a comprehensive understanding of the visual content, which is essential for generating accurate and meaningful Q&A pairs.
@@ 2.2 Question Answer Pair Generation
Next, we generate corresponding Q&A pairs for the extracted image and text information, respectively. After reorganizing them through DSG, they can be directly input into the model to detect the level of hallucination.
We use two kinds of questions for hallucination evaluation. The first type is Yes-No questions, which check object existence, such as “Is there any {objk}?" and object relations like “Is there a {obj1} near the {objk}?". These questions help determine whether specific objects and their relationships are present within an image. Additionally, Yes-No questions assess attributes by asking about features like color, size, or location.
The second type is Wh-Questions, which add diversity to the evaluation by incorporating interrogative words such as “what”, “who", “which”, “where”, and “how many”. These questions elicit more detailed, free-form responses, typically no longer than three words. For example, “What color is the {objk}?" or “Which object is near the {objk}?" help assess the finer details about objects and their relationships. Unlike traditional hallucination evaluations that primarily rely on Yes-No questions, our approach includes both types to provide a more comprehensive assessment.
We introduce Negative Questions for both Yes-No and Wh-Questions. These questions are created by replacing real objects, attributes (e.g., color, size), and relations in the original Q&A pairs with non-existent ones. For example, "Is there a car here?" becomes "Is there a plane here?" using a randomly selected object from a pre-constructed set. To ensure accurate evaluation, we avoid including objects present in the current image in the candidate set. These questions are answered with negative pronouns like "none," "nobody," or "nowhere."
@@ 2.3 Davidson Scene Graph
To model the dependency between objects, attributions, and relations accurately and improve the reliability of hallucination evaluation, we introduce the Davidsonian Scene Graph (DSG) Cho et al. (2023) mechanism. The DSG schematic diagram can refer to Figures 1 and 8. The DSG can be understood as a post-processing step for the Q&A pairs. After obtaining all the Q&A pairs, we organize them into multiple tree-like structures, where each Q&A pair serves as a node. According to the structure of the tree-like structures, each node is either a root node or a leaf node. Specifically, the entire process is divided into three steps. In step 1, we set the question about the existence of a certain object as the root node. In step 2, we set all questions related to the object of the root node, such as those about its attributes and relations, as corresponding leaf nodes. Finally, in step 3, determine whether the root node question is answered correctly; if not, there is no need to judge the questions at the leaf nodes, and we directly determine that all questions on the tree are answered incorrectly. For instance, after step 1 and step 2, we obtain a list of questions such as $L^{Q}=\{Q_{1}:Independent,Q_{2}:Depends\ on\ Q_{1}\}$. Before determining if the answer to $Q_{2}$ is correct, we first assess $Q_{1}$, which concerns the accuracy related to the root node. If the question about the existence of an object, which is at the root node, is answered incorrectly, we consider that all other related questions must be hallucinatory.
@@ 3 Experiments
Table 2: The number of questions generated by FIHA from various datasets.
Source From Image From Caption
MSCOCO 25,699 13,007
Foggy 7,232 2,801
Visual Genome 1,566 476
Table 3: Evaluation results of LVLMs on questions generated from images and captions using FIHA. The upper part is from the MSCOCO dataset and the bottom part is from the Foggy dataset. F1 (Gen) is the BERTScore value calculated from the standard answers and model outputs for all Wh-Questions. For more details, please refer to the explanation of Metrics in Section 3.1.
@@
mPLUG-Owl 42.1 70.2 61.4 43.7 15.2 31.4 61.6 55.5 31.2 11.4
MiniGPT-4 23.5 27.5 22.2 22.1 21.6 15.9 25.7 28.8 14.2 18.4
MultiModal-GPT 59.1 46.4 47.1 46.6 16.1 23.8 39.6 45.7 22.1 10.8
LLaVA-1.5-7B 77.8 77.0 65.9 67.7 21.4 50.7 64.9 67.5 50.5 13.7
LLaVA-1.5-13B 78.9 80.9 66.4 68.3 20.9 47.6 64.2 65.5 48.5 13.8
InstructBLIP 84.7 83.3 78.6 80.4 21.8 65.7 69.5 77.4 64.2 14.1
GPT-4V 87.2 81.4 86.3 85.5 25.2 70.3 71.5 75.8 69.3 22.7
@@
mPLUG-Owl 64.8 60.2 51.1 42.7 18.6 29.5 58.9 51.6 25.6 29.3
MiniGPT-4 30.1 30.2 27.6 28.1 9.4 23.4 34.4 37.8 23.0 11.6
MultiModal-GPT 50.2 48.7 46.1 45.8 17.6 28.1 43.9 47.9 25.4 24.5
LLaVA-1.5-7B 67.7 68.4 56.2 52.9 19.7 29.1 50.0 49.2 25.8 27.5
LLaVA-1.5-13B 68.1 71.5 56.1 52.3 18.8 28.9 49.2 49.8 25.5 27.7
InstructBLIP 70.9 75.6 60.2 58.8 20.3 32.8 58.3 53.2 30.5 29.2
GPT-4V 76.3 70.1 64.6 66.0 16.2 33.7 53.3 51.7 32.1 21.7
@@
mPLUG-Owl 41.8 68.9 60.9 43.3 16.4 44.6 71.7 51.8 33.8 24.1
MiniGPT-4 22.9 26.8 22.0 21.8 22.3 15.9 25.7 28.8 14.2 16.8
MultiModal-GPT 58.8 46.1 46.9 46.3 17.2 65.3 65.2 62.2 61.7 21.7
LLaVA-1.5-7B 77.7 77.2 61.1 67.9 20.6 56.4 73.8 62.0 52.6 20.1
LLaVA-1.5-13B 79.0 81.2 66.7 68.6 20.9 74.3 79.6 68.8 69.2 20.2
InstructBLIP 84.5 83.7 79.0 80.7 22.4 67.7 78.4 71.9 66.7 26.8
GPT-4V 87.0 81.2 86.0 85.3 23.7 84.2 78.9 84.1 82.2 26.0